# Normality-Guided Distributional Reinforcement Learning for Continuous Control

### v3

@@ 6 Conclusion
We have presented a distributional reinforcement learning (DRL) method in that output quantiles approximate the value distribution as a normal distribution under the mild assumption for Markov Chain Central Limit Theorem Jones (2004). Existing actor-critic algorithms that assess uncertainty typically employ multiple value functions (ensemble) to estimate variance. Given the presence of states with relatively high variance, we propose an alternative approach to measure uncertainty. Rather than relying on the variance, we evaluate how much the predicted quantiles are close to a normal.
@@ Acknowledgement
The authors are thankful for the generous support of computational resources by the Ohio Supercomputer Center
@@ Appendix A Hyperparameters
Deep reinforcement learning algorithms pose a significant challenge for evaluation due to the inherent instability and stochasticity during training. To ensure a fair comparison with other algorithms, we attempted to maintain a consistent set of hyperparameters as much as possible. We follow Spinning Up Achiam [2018] default settings for learning rate, optimizer, etc. All policies have a two-layer tanh network with 64 x 32 units, and all of the value function and the distributional value function has a two-layer ReLU network with 64 x 64 units or 128 x 128 units for all environments.
Although Lee et al. Lee et al. [2020] suggest the formula to compute the uncertainty weight $w(s)=\sigma(-E*T)+w_{\rm{min}}$, it seems that finding the reasonable value for $T$ is not intuitive for various tasks. (Note that $E$ is computed as the normality error of the distributional value function in our method and $E$ is replaced with the standard deviation of the value functions in Lee et al. Lee et al. [2020]). We instead set a target uncertainty weight and perform a parametric search by adjusting $T$ to achieve that target weight (Appendix B). This process substitutes $T$ with the target weight $w_{\rm{tar}}$ as a hyperparameter. $w_{\rm{tar}}$ is chosen from $w_{\rm{tar}}\in\{0.85,0.9\}$, and $w_{\rm{min}}$ is chosen from $w_{\rm{min}}\in\{0.4,0.5,0.6\}$.
MC-CLT has the number of quantile output nodes of the distributional value function as a hyperparameter. Our observations indicate that having the number of quantile output nodes greater than or equal to 8 yields satisfactory performance; thus, we fix this number across all environments. Also, the ensemble-based method introduces the ensemble size as a hyperparameter. As Lee et al. Lee et al. [2020] and Mai et al. Mai et al. [2022] demonstrate that ensemble size $=5$ shows the sufficient results, we set the ensemble size to $5$ for these models.
@@ Appendix B Algorithms In Detail
1:  Input: target weight $w_{\rm{tar}}$,
2:      minimum weight $w_{\rm{min}}$, error set $E$
3:  Initialize $left=0,right=2^{12}$
4:  while $left\leq right$
5:     $T=(left+right)/2$
6:     $W=2(1-w_{\rm{min}})\sigma(-E*T)+w_{\rm{min}}$
7:     Compute the mean of uncertainty weight
8:         $w_{\rm{avg}}=\rm{Avg}(W)$
9:     if $w_{\rm{avg}}$ is in the range $[w_{\rm{tar}}-\epsilon,w_{\rm{tar}}+\epsilon]$
10:        Use $T$ to compute the uncertainty weight
11:        break
12:     elif $w_{\rm{avg}}\geq w_{\rm{tar}}+\epsilon$
13:        $left=T$
14:     else
15:        $right=T$
16:  end while
@@ Appendix C Variance Network
In practice, the variance network Kendall and Gal [2017] is trained to predict log variance $\rm{ln}\sigma^{2}_{\psi}$ instead of predicting variance directly. This avoids a division by zero or a negative value in the log term in Equation (6). However, after predicting log variance, it is necessary to retrieve the actual variance value through the exponential operation, i.e., $\sigma^{2}_{\psi}=\rm{exp}(\rm{ln}\sigma^{2}_{\psi})$.
@@ Appendix D Standard Deviation Change
Certain states display relatively higher variances compared to others. For instance, in fragile control tasks like Walker2d-v2, the states close to the done state typically exhibit higher variances. This is because once the agent overcomes the done states, several more successful actions are likely to be generated, increasing the sum of rewards (return). Additionally, the closer the state is to the initial timestep, the higher the variance. This correlation is apparent when comparing the changes in return from the initial and intermediate timesteps.
The graphs (Figure 2) shown below have been obtained by applying exponential smoothing to the standard deviation of the trained value functions (the ensemble-based method), Notably, the standard deviation is observed to decrease over the timestep.
@@ Appendix E Normality Discussion
The question arises: What happens when we use the uncertainty weights to update the policy in tasks deviating from the normality condition? In this case, we argue that our method becomes a special case of Exploration by Random Network Distillation (RND) Burda et al. [2018], in which a random neural network is provided, and another neural network is fitted to this random network. While the model is fitted, an exploration bonus is computed based on the differences between the two models. Differing from RND’s approach of providing a random neural network, we provide a normal quantile target, expecting that $V^{D}$ will also produce quantile bars indicative of a normal distribution. We then measure how closely $V^{D}$’s predictions mirror the shape of a normal distribution.
Both approaches rely on the premise that the difference diminishes with sufficient learning. The distinction between RND and our method is that RND utilizes uncertainty as a bonus added to the reward, seeking novelty, whereas our approach is a risk-averse strategy by reducing the impact of gradients from high-uncertainty samples. These different search strategies can be chosen depending on the problem. Given that the reward function in control tasks, such as MuJoCo, is highly engineered, being risk-averse is more effective than seeking novelty, as the feedback from the reward function is enough unlike sparse reward settings.
@@ Appendix G Potential Research Directions
The use of our method does not exclude the possibility of utilizing the ensemble technique. In fact, by combining our method with the ensemble, it becomes possible to leverage multiple distributional value functions. However, simply averaging the uncertainties from each function has shown no performance improvement compared to our method alone. Besides weighting policy updates, exploring alternative uses of uncertainties from each distributional value function would be an intriguing avenue for future research. For instance, in methods like Soft Actor-Critic (SAC) Haarnoja et al. [2018] or Twin Delayed DDPG (TD3) Fujimoto et al. [2018], where the minimum value is selected among two Q functions, the Q function with lower uncertainty could be chosen to utilize a more accurate value because our uncertainty weight represents the accuracy of the network.

### v4

@@ 6 Conclusion
We have presented a distributional reinforcement learning (DRL) method in that output quantiles approximate the value distribution as a normal distribution under the mild assumption for Markov Chain Central Limit Theorem (Jones, 2004). Existing actor-critic algorithms that assess uncertainty typically employ multiple value functions (ensemble) to estimate variance. Given the presence of states with relatively high variance, we propose an alternative approach to measure uncertainty. Rather than relying on the variance, we evaluate how much the predicted quantiles are close to a normal.
@@ Acknowledgement
The authors would like to thank the Ohio Supercomputer Center (Center, 1987) for providing the computational resources used in this research.
@@ Appendix A Hyperparameters
Deep reinforcement learning algorithms pose a significant challenge for evaluation due to the inherent instability and stochasticity during training. To ensure a fair comparison with other algorithms, we attempted to maintain a consistent set of hyperparameters as much as possible. We follow Spinning Up (Achiam, 2018) default settings for learning rate, optimizer, etc. All policies have a two-layer tanh network with 64 x 32 units, and all of the value function and the distributional value function has a two-layer ReLU network with 64 x 64 units or 128 x 128 units for all environments.
Although Lee et al. (2020) suggest the formula to compute the uncertainty weight $w(s)=\sigma(-E*T)+w_{\rm{min}}$, it seems that finding the reasonable value for $T$ is not intuitive for various tasks. (Note that $E$ is computed as the normality error of the distributional value function in our method and $E$ is replaced with the standard deviation of the value functions in Lee et al. (2020)). We instead set a target uncertainty weight and perform a parametric search by adjusting $T$ to achieve that target weight (Appendix B). This process substitutes $T$ with the target weight $w_{\rm{tar}}$ as a hyperparameter. $w_{\rm{tar}}$ is chosen from $w_{\rm{tar}}\in\{0.85,0.9\}$, and $w_{\rm{min}}$ is chosen from $w_{\rm{min}}\in\{0.4,0.5,0.6\}$.
MC-CLT has the number of quantile output nodes of the distributional value function as a hyperparameter. We examine how performance is affected by the number of quantile bars, with values such as $N=4,8,12,$ and $20$. We observe satisfactory performance when the number of quantile bars is set to $N\geq 8$, so we fix the number of quantiles at 8 for all experiments. Also, the ensemble-based method introduces the ensemble size as a hyperparameter. As Lee et al. (2020) and Mai et al. (2022) demonstrate that ensemble size $=5$ shows the sufficient results, we set the ensemble size to $5$ for these models.
@@ Appendix B Algorithms In Detail
1:Input: target weight $w_{\rm{tar}}$, minimum weight $w_{\rm{min}}$, error set $E$
2:Initialize $left=0$, $right=2^{12}$
3:while $left\leq right$ do
4:     $T=(left+right)/2$
5:     $W=2(1-w_{\rm{min}})\sigma(-E*T)+w_{\rm{min}}$
6:     Compute the mean of uncertainty weight $w_{\rm{avg}}=\mathrm{Avg}(W)$
7:     if $w_{\rm{avg}}$ is in the range $[w_{\rm{tar}}-\epsilon,w_{\rm{tar}}+\epsilon]$ then
8:         Use $T$ to compute the uncertainty weight
9:         break
10:     else if $w_{\rm{avg}}\geq w_{\rm{tar}}+\epsilon$ then
11:         $left=T$
12:     else
13:         $right=T$
14:     end if
15:end while
@@ Appendix C Variance Network
In practice, the variance network (Kendall & Gal, 2017) is trained to predict log variance $\rm{ln}\sigma^{2}_{\psi}$ instead of predicting variance directly. This avoids a division by zero or a negative value in the log term in Equation (6). However, after predicting log variance, it is necessary to retrieve the actual variance value through the exponential operation, i.e., $\sigma^{2}_{\psi}=\rm{exp}(\rm{ln}\sigma^{2}_{\psi})$.
@@ Appendix D Standard Deviation Change
Certain states display relatively higher variances compared to others. For instance, in fragile control tasks like Walker2d-v2, the states close to the terminal state typically exhibit higher variances. This is because once the agent overcomes the terminal states, several more successful actions are likely to be generated, increasing the sum of rewards (return). Additionally, the closer the state is to the initial timestep, the higher the variance. This correlation is apparent when comparing the changes in return from the initial and intermediate timesteps.
The graphs (Figure 4) have been obtained by applying exponential smoothing to the standard deviation of the trained value functions (the ensemble-based method), Notably, the standard deviation is observed to decrease over the timestep.
@@ Appendix E Normality Discussion
The question arises: What happens when we use the uncertainty weights to update the policy in tasks deviating from the normality condition? In this case, we argue that our method becomes a special case of Exploration by Random Network Distillation (RND) (Burda et al., 2018), in which a random neural network is provided, and another neural network is fitted to this random network. While the model is fitted, an exploration bonus is computed based on the differences between the two models. Differing from RND’s approach of providing a random neural network, we provide a normal quantile target, expecting that $V^{D}$ will also produce quantile bars indicative of a normal distribution. We then measure how closely $V^{D}$’s predictions mirror the shape of a normal distribution.
Both approaches rely on the premise that the difference diminishes with sufficient learning. The distinction between RND and our method is that RND utilizes uncertainty as a bonus added to the reward, seeking novelty, whereas our approach is a risk-averse strategy by reducing the impact of gradients from high-uncertainty samples. These different search strategies can be chosen depending on the problem. Given that the reward function in control tasks, such as MuJoCo, is highly engineered, being risk-averse is more effective than seeking novelty, as the feedback from the reward function is sufficient, unlike sparse reward settings. We actually conduct experiments to check whether encouraging the agent to visit high-uncertainty states is beneficial for our environments (Table 3), and find that seeking exploration does not outperform in environments with highly engineered rewards.
@@ Appendix G Potential Research Directions
The use of our method does not exclude the possibility of utilizing the ensemble technique. In fact, by combining our method with the ensemble, it becomes possible to leverage multiple distributional value functions. However, simply averaging the uncertainties from each function has shown no performance improvement compared to our method alone. Besides weighting policy updates, exploring alternative uses of uncertainties from each distributional value function would be an intriguing avenue for future research. For instance, in methods like Soft Actor-Critic (SAC) (Haarnoja et al., 2018) or Twin Delayed DDPG (TD3) (Fujimoto et al., 2018), where the minimum value is selected among two Q functions, the Q function with lower uncertainty could be chosen to utilize a more accurate value because our uncertainty weight represents the accuracy of the network.
Table 3: Seeking Accuracy vs. Seeking Exploration Table: We compute the uncertainty weight (our method) based on Equation 9. However, this raises concerns regarding exploration. To address this, we conduct an experiment using Equation (2) to compute the exploration weight, ensuring that states with high uncertainty are visited more frequently. Each entry represents an average return and a standard error, derived from 30 different settings, with each setting running for 100 episodes.
BipedalWalker Hopper HalfCheetah Ant
Exploration TRPO $34\pm 3.3$ $2328\pm 34$ $2459\pm 32$ $1663\pm 16$
MC-CLT TRPO (ours) $\mathbf{195\pm 2.6}$ $\mathbf{2927\pm 27}$ $\mathbf{3172\pm 43}$ $\mathbf{3329\pm 20}$
Swimmer Walker2d InvertedDoublePendulum LunarLander(C)
Exploration TRPO $36\pm 2.2$ $1086\pm 33$ $7628\pm 112$ $179\pm 4.4$
MC-CLT TRPO (ours) $118\pm 0.2$ $\mathbf{2904\pm 23}$ $\mathbf{8404\pm 93}$ $\mathbf{258\pm 1.8}$
BipedalWalker Hopper HalfCheetah Ant
Exploration PPO $70\pm 3.6$ $2727\pm 28$ $2756\pm 37$ $1315\pm 10$
MC-CLT PPO (ours) $\mathbf{236\pm 2.3}$ $\mathbf{3230\pm 22}$ $\mathbf{3110\pm 39}$ $\mathbf{1900\pm 13}$
Swimmer Walker2d InvertedDoublePendulum LunarLander(C)
Exploration PPO $44\pm 0.8$ $916\pm 24$ $9027\pm 54$ $230\pm 2.8$
MC-CLT PPO (ours) $\mathbf{122\pm 0.1}$ $\mathbf{2716\pm 33}$ $\mathbf{9225\pm 36}$ $\mathbf{278\pm 1.2}$